Find the sum of: ( (d) – 50, – 200 and 300
asdioghrsiounfinihsdriuohbrdi

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(-50) + (-200) + (+300) = (-50) + (-200) + (+200) + (+100)
=(-50) + 0 + (+100)
= (-50) + (+100)
= (-50) + (+50) + (+50) = 0 + (+50)
= 50

Thanks